OpenStudy (nathanjhw):

Suppose f(x)=4x^5 - (1/x^4). Find the slope of the tangent line to the graph of y = f^(–1)(x) at (3, 1)

OpenStudy (nathanjhw):

1/24 1/16 00062 1620 972

OpenStudy (nathanjhw):

@iambatman

OpenStudy (nathanjhw):

@Zale101

OpenStudy (perl):

there is a theorem for finding derivative of inverse

OpenStudy (perl):

$$ \Large (f^{-1}(x))' = \frac{1}{f' (f^{-1}(x))} $$

OpenStudy (perl):

do you see how to use it ?

OpenStudy (nathanjhw):

When I took the inverse of the function and then derived it I got: (20x^9 + 4)/ x^5

OpenStudy (perl):

is this the original function $$ \Large f(x) = 4x^5 - \frac{1}{x^4} $$

OpenStudy (nathanjhw):

Yes.

OpenStudy (nathanjhw):

Yes.

OpenStudy (perl):

and can you show me how you got the inverse, did you switch x and y ?

OpenStudy (nathanjhw):

Oh I think I derived the original function and not the inverse.

OpenStudy (perl):

it is actually not necessary to find the inverse function explicitly , because of this theorem $$ \Large{ (f^{-1}(x))' = \frac{1}{f' (f^{-1}(x))} \\ But...\\ f(x) =4x^5 - \frac{1}{x^4} = 4x^5 - x^{-4} \\ f ' (x) = 4(5)x^4 - (-4)x^{-5} = 20x^4 + \frac{4}{x^5} \\ \therefore \\ (f^{-1}(3))' = \frac{1}{f' (f^{-1}(3))} = \frac{1}{f' (1)} = \frac{1}{20\cdot 1 + \frac{4}{1}} } $$

OpenStudy (nathanjhw):

Wow, using the formula makes it so much easier. Thank you so much for the help!